The 500-N weight is supported by two cables, the cable forces being F1 and F2. Knowing that the resultant of F1 and F2 is a force of magnitude 500 N acting in the y-direction, determine F1 and F2.


1
Expert's answer
2021-06-01T14:48:44-0400

Gives

"F=500N"

Assume diagram



Force horizontal components

"\Sigma F_H=0"

"-T_1cos 35\u00b0+T_2cos50\u00b0=0"

"T_2=\frac{T_1cos35}{cos50}"

Vertical components

"\Sigma F_v=0"

"T_1sin35\u00b0+T_2sin50\u00b0-500=0"


Put T2 value


"T_1sin35°+\frac{T_1cos35}{cos50}\times sin50°-500°=0"



"T_1=322.631477N"


"T_2=T_1\frac{cos35°}{cos50°}"

"T_2=411.14N"